# System Identification of Neural Systems: Going Beyond Images to Modelling Dynamics

### v1

@@ 3.3 Candidate models
Here we describe the candidate models that are used in our experiments. We choose to run our experiments on more than 30 source models that are listed in Table 1, along with their model family and configurations. Video understanding models include C2D Li et al. (2019), CSN Tran et al. (2019), I3D Carreira & Zisserman (2017), R(2+1)D Tran et al. (2018), SlowFast and the Slow branch (3D ResNet-50) Feichtenhofer et al. (2019), X3D Feichtenhofer (2020), MViT Fan et al. (2021) and TimeSformer Bertasius et al. (2021). Self-supervised video understanding models, stMAE Feichtenhofer et al. (2022) and OmniMAE Girdhar et al. (2023) are used as well. Single image models include ResNets He et al. (2016), ViTs Dosovitskiy et al. (2021) and the self-supervised DINO Caron et al. (2021) and MAE He et al. (2022). Families of models are categorized based on the input, supervision and architecture type as discussed earlier. Since our focus is on video understanding, most of our models are from the video understanding literature. We also detail the number of layers and which layers are sampled for each deep network used in our experiments in Table 2. Note that for transformer-based architectures instead of using all layers to be selected we rather group layers into blocks of four layers for efficiency reasons and we noticed it gave better results than learning the regression with all input layers at once. Also note that OmniMAE finetuned is trained for the action recognition task on SSV2 dataset Goyal et al. (2017).
@@ 4.1 Implementation details
In this subsection, we describe our experiment design and details of our implementation for both the simulated and real sets of experiments. Across both, there are common hyper-parameters that we use. In the case of both video and image understanding, we sample a clip from the input video to extract features for that clip. In image understanding models, we extract features per frame and combine the output features from the clip. The input clips to all models are constructed based on a sampling rate that corresponds to the sampling rate used during its training for video understanding models. As for image understanding models we use the default sampling rate eight. The clip length is computed based on the sampling rate and the input video length and changes according to the video length.
@@
Before training the regressor, a hyperparameter tuning is conducted using two-fold cross-validation on the training set of the first subject, following previous work Zhou et al. (2022). The two main hyper-parameters we tune are the $\beta_{1},\beta_{2}$, using grid search $\beta_{1}\in\{0.1,1,10\}$ and $\beta_{2}\in\{1,10,100\}$. Moreover, an early-stopping strategy is employed through the hyperparameter tuning and training phases. The main metric used throughout the experiments is the average Pearson’s correlation coefficient across all voxels within a specific region of interest in the brain. All results are averaged over the subjects. We conduct experiments on four folds and report the average and standard deviation. We report the statistical significance across families of models using Welch’s t-test.
The real setup is conducted on the Mini-Algonouts2021 dataset Cichy et al. (2019) of fMRI responses recorded from ten subjects who watched three repetitions of 1,000 video stimuli with an average duration of three seconds. In our analysis, we averaged the fMRI responses across the video repetitions to represent each single video with one activation value at each voxel. In each fold of the four folds we used in our experiments, the 1,000 videos are split into training and testing sets as 90% and 10%, respectively. Each of our candidate models’ features are used to regress on the fMRI data with the regularized regression that allows for re-weighing the features from the different layers for each region in the brain. In the simulated setup, we conduct it on three target models and follow the same setup for the real environment with the video stimuli used as input to the target deep network.
4.2 Can we perform system identification with respect to the dynamics modelling?
We first investigate the question of whether system identification of single image vs. video understanding models is possible. Towards that end, we study two families of models in a simulated environment with five single image and 13 video understanding models as source models, and the target models defined in Section 3.3. Figure 1 shows that for MViT and ViT target models, the correct family of models (either image or video family) is better able to represent each. Specifically, we see the video understanding models for MViT and the single image one for ViT show higher regression scores. In the case of MViT, all layers show statistically significant results between image and video families of models. As an example, video understanding models achieved an average score of 0.58 while image understanding ones achieved an average score of 0.49 in representing the first block (i.e., third layer) in MViT (presented as B1 in the figure). As for ViT, it shows statistically significant results across all layers except the last one. Note that we include in the video understanding family, models that are trained on three different datasets which are Kinetics, Charades, and Something-Something v2 to ensure the results are not dependent on a certain training dataset.
When looking at I3D ResNet-50, we notice that the difference between the two families of models is statistically insignificant across most layers, especially the early ones. Our interpretation of these results is tied to one of the recent interpretability studies of video understanding models Kowal et al. (2022). The aforementioned study showed that I3D was the weakest in modelling dynamics with respect to other video understanding models. Hence, it might be the reason behind the insignificant difference between image and video-based models in the early layers when I3D is the target.
In summary, we have demonstrated that with MViT and ViT there is a strong indication that system identification can be attained to a certain level with most of the layers showing statistical significance. Additional simulated experiments are provided in Appendix A.1. Since we can identify the ground truth modelling scheme in terms of image vs. video understanding in the simulated setup, in the following section we investigate the same but for the unknown biological neural system.
@@ 4.2 Can we perform system identification with respect to the dynamics modelling?
Figure 1: Simulated experiments showing regression scores as Pearson’s correlation coefficient of image (blue) vs. video (red) model families on three target models; (a) MViT-B, (b) ViT-B, (c) I3D ResNet-50. Statistical significance is shown at the bottom as ‘ns’ not significant, ‘*,**,****,**,**** , * * , * * *’ significant with p-values $<0.05,0.01,0.001$, resp.
4.3 How does the human visual cortex compare to deep networks when taking dynamics modelling into consideration?
Since we have established the feasibility of identifying the target image vs. video understanding models to an extent with regression scores, it brings the question of how can we use this information to identify the underlying mechanisms in biological neural systems. The results for the real experiments, where the target model is the biological neural system, are shown in Figure 2. We conduct three comparisons; a comparison of single image vs. video understanding families of models, a comparison of convolutional-based vs. transformer-based models, and a comparison of fully-supervised vs. self-supervised models. Figure 2 (a) clearly demonstrates that across most brain regions, video understanding models have better capability to model the visual cortex responses than single image architectures. As an example, video understanding models have an average score of 0.28 in predicting V1 region responses while image understanding ones have an average score of 0.24. This result questions the previous large-scale studies when comparing deep networks to biological neural systems, as they have missed a crucial aspect that should be considered which is modelling dynamics. Moreover, our results are statistically significant across all regions except the PPA which is concerned with scenes and might be giving less weight to modelling dynamics.
@@
Figure 2 (b) shows the comparison between transformer-based and convolutional-based models. It shows that convolutional models have higher regression scores across early-mid regions in the visual cortex with relatively high statistical significance. Interestingly, it has been noted in recent works that transformers lack the ability to capture high-frequency components Bai et al. (2022). On the other hand, early layers in convolutional models tend to capture high-frequency components by detecting oriented gradients. This also might relate to empirical results that demonstrated vision transformers with shallow convolutional stem (i.e., three convolutional layers) perform better than the ones that directly take patches as input Xiao et al. (2021). As such, brain modelling in the early and mid regions of the visual cortex relates better to convolutional-based models, as they are concerned with high-frequency components. It also brings insight into the design of transformer-based models to encourage the adoption of shallow convolutional stem widely in transformers, beyond the patchifying stem that renders vision transformers as less capable of capturing high-frequency components.
Surprisingly, Figure 2 (c) shows that fully-supervised models are better able to model the early regions than self-supervised models. It also shows that the differences between fully and self-supervised are statistically significant across the four early regions of the visual cortex (V1, V2, V3, and V4). However, the results are not significant in the later regions (LOC, EBA, FFA, STS, and PPA). Additionally, in Appendix A.2, we studied the model subfamilies while focusing only on the video understanding models and excluding the single image models. In the following sections, we will show a more fine-grained analysis of such a result to confirm it.
@@
Figure 2: Real experiments showing regression scores as Pearson’s correlation coefficient of model families on brain fMRI data. (a) Comparison of image vs. video understanding models, (b) comparison of convolutional vs. transformer-based models and (c) comparison of fully supervised vs. self-supervised models. Statistical significance is shown in the bottom as ‘ns’ not significant, ‘*,**,****,**,**** , * * , * * *’ significant with p-values $<0.05,0.01,0.001$, resp.
@@ 4.4 Fine-grained analysis
Figure 3: Fine-grained analysis of the video understanding models across the nine regions of the human visual cortex showing the Pearson’s correlation coefficient as the regression scores. (a) Comparison between single stream and two stream SlowFast architectures. (b) Comparison between OmniMAE pre-trained in a self-supervised manner, TimeSformer and OmniMAE fine-tuned to a downstream task with full supervision. All models are based on ViT-B and trained on SSV2. (c) Comparison between six video understanding models. Statistical significance is shown on top of bar pairs as ‘ns’ not significant, ‘*,**,****,**,**** , * * , * * *’ significant with p-values $<0.05,0.01,0.001$, resp.

### v2

@@ 3.2 Candidate models
Here we describe the candidate models that are used in our experiments. We choose to run our experiments on more than 35 source models, listed in Table 1, with their model family and configurations. Video understanding models include C2D Li et al. (2019), CSN Tran et al. (2019), I3D Carreira & Zisserman (2017), R(2+1)D Tran et al. (2018), SlowFast, the Slow branch (3D ResNet-50) Feichtenhofer et al. (2019), X3D Feichtenhofer (2020), MViT Fan et al. (2021b) and TimeSformer Bertasius et al. (2021). Self-supervised video understanding models, stMAE Feichtenhofer et al. (2022) and OmniMAE Girdhar et al. (2023) are used as well. We mainly focus on the models that showed state of the art performance in action recognition. Single image understanding models include ResNets He et al. (2016), ViTs Dosovitskiy et al. (2021) and the self-supervised DINO Caron et al. (2021) and MAE He et al. (2022). Families of models are categorized based on the input, supervision and architecture type as discussed earlier. We detail the number of layers and which layers are sampled for each deep network used in Appendix A.
Figure 1: Architecture of our dynamics based neural encoding with inter-intra region connectivity priors. Our fully integrated model learns the connectivity from all regions voxels (inter-region) and all voxels in the same region (intra-region). We only show one target region, V1, as an illustration where we use the same mechanism across all regions. Note the arrows thickness, in our visual cortex regions illustration, indicate the degree of connectivity corresponding to the computed in Fig. 5(c).
3.3 Encoding technique
Inspired by the recent work Zhou et al. (2022), we use a layer-weighted region of interest encoding that takes the hierarchical nature of deep networks into consideration. Initially, we sample the frames from a video to obtain the input clip. For image understanding models we extract features per frame, while for video ones we extract features from the entire clip at once. Then, we pre-process the input features from the different layers of a candidate model through averaging the features on the temporal dimension. This is followed by performing sparse random projection Li et al. (2006) for dimensionality reduction and computational efficiency reasons. Assume input features for layer, $l$, after dimensionality reduction as, $X_{l}\in\mathbb{R}^{C\times 1}$, with $C$ features. We learn the weights of one fully connected layer to provide the predictions of the voxels of one region of interest in the visual cortex as, $\hat{Y}_{l}=W_{l}X_{l}$. Where $W_{l}\in\mathbb{R}^{N\times C}$, $\hat{Y}\in\mathbb{R}^{N\times 1}$ and $N$ is the number of voxels in the region of interest. Instead of a simple ridge regression, we learn a weighted sum of the predictions of all layers and use the following loss to train our regression model,
3.4 Inter-intra region connectivity priors
In this section, we present a novel encoding scheme on top of the best-performing video understanding models by fully integrating the neural encoding with inter- and intra-region voxel connectivity priors. An overview of our architecture is shown in Fig. 1. The input video stimuli goes through the source video understanding model to extract multiple layers features as described in Sec. 3.3, followed by the connectivity module which takes the concatenated voxels of the nine visual regions as input. This module consists of four layers: two fully-connected coupled with L2 regularization and two dropout ones. Finally, the output of the model is the predicted voxels of a single visual region. We train our model in a two-stage fashion, where we train the standard neural encoding scheme without connectivity, followed by training the connectivity module. In the training phase, the main target is to learn the connectivity between the voxels of all the regions and the target region including intra-connectivity between the voxels of the target region itself and the inter-connectivity between voxels of the target region and the other visual regions. Accordingly, in the training phase, the input to the model is the groundtruth voxel activations. However, in the inference phase, the input to the model is the predicted voxel activations.
@@ 4.1 Implementation details
In this subsection, we describe our experiment design and implementation details. The input clips to all models are constructed based on a sampling rate that corresponds to the sampling rate used during its training for video understanding models. As for image understanding models we use the default sampling rate of eight. The clip length is computed based on the sampling rate and the input video length and changes according to the video length.
@@
Before training the regressor, a hyperparameter tuning is conducted using two-fold cross-validation on the training set of the first subject, following previous work Zhou et al. (2022). The two main hyper-parameters we tune are the $\beta_{1},\beta_{2}$, using grid search $\beta_{1}\in\{0.1,1,10\}$ and $\beta_{2}\in\{1,10,100\}$. Moreover, an early-stopping strategy is employed through the hyperparameter tuning and training phases. The main metric used throughout the experiments is the average Pearson’s correlation coefficient across all voxels within a specific region in the brain. All results are averaged over the subjects. We conduct experiments on four folds and report the average and standard deviation. We report the statistical significance across families of models using Welch’s t-test. The real setup, in which the human visual cortex is the target, is conducted on the Mini-Algonouts 2021 dataset Cichy et al. (2019), additional results on BOLD moments dataset is provided in Appendix B.1. In each fold of the four folds we used in our experiments, the 1,000 videos are split into training and testing sets as 90% and 10%, respectively. In the simulated setup, we follow the same setup for the real environment. All the previous models are trained on A6000 GPU with training span of one day average per regression model.
@@
Figure 2: Simulated experiments showing regression scores as Pearson’s correlation coefficient of image (blue) vs. video (red) model families on four target models; (a) MViT-B, (b) ViT-B, (c) ResNet-50, (d) I3D ResNet-50. We show the regression on the target network output features from their respective blocks, B0-7. Statistical significance is shown at the bottom as ‘ns’ not significant, ‘,,*,**,***∗ , ∗ ∗ , ∗ ∗ ∗’ significant with p-values $<0.05,0.01,0.001$, resp. It shows higher regression scores for the model family corresponding to the target network, especially in MViT and ViT.
4.2 Can we perform system identification with respect to the dynamics encoding?
We first investigate the question of whether system identification of single image vs. video understanding models is possible. Figure 2 shows that for MViT and ViT target models, the correct family of models is better able to represent each. Specifically, we observe higher regression scores of the video understanding models for MViT and the single image one for ViT. Note that we include video understanding models that are trained on three different datasets which are Kinetics, Charades, and Something-Something v2 to ensure the results are not dependent on a certain training dataset.
When inspecting ResNet-50 as a target, we observe the mean of image understanding models is higher, but is only statistically significant in three layers. When looking at I3D target model, we notice the mean of video understanding models is higher and is statistically significant except in early layers. In summary, we have demonstrated that system identification can be attained to a certain level with most of the layers showing statistical significance. Additional simulated experiments are provided in Appendix B.2. In the following section we investigate the same but for the unknown biological neural system.
4.3 How do deep video understanding models families compare to the human visual cortex?
@@
In this section, we focus on the real experiments, where the target model is the biological neural system, to understand the underlying mechanisms of the visual cortex. We conduct three comparisons; single image vs. video understanding families of models, convolutional-based vs. transformer-based models, and fully-supervised vs. self-supervised models. Figure 3(a) clearly demonstrates that across most brain regions, video understanding models have better capability to model the visual cortex responses than single image architectures. We believe the reason behind this is that the brain is encoding the dynamics occurring in a video similar to the encoded dynamics in video understanding models more than what occurs in single image understanding models. Figure 3(b) shows the comparison between transformer-based and convolutional-based models. It shows that convolutional models have higher regression scores across early-mid regions in the visual cortex with relatively high statistical significance. This difference decrease as we go to higher level regions until it becomes insignificant. Interestingly, it has been noted in recent works that transformers lack the ability to capture high-frequency components Bai et al. (2022). On the other hand, early layers in convolutional models tend to capture high-frequency components by detecting oriented gradients. This also might relate to empirical results that demonstrated vision transformers with shallow convolutional stem (i.e., three convolutional layers) perform better than the ones that directly take patches as input Xiao et al. (2021). As such, brain modelling in the early and mid regions of the visual cortex relates better to convolutional-based models, since they better capture high-frequency components. Nonetheless, we notice the best model in the transformer based family, MViT, tends to behave similar to convolutional ones in early-mid regions unlike other transformer based architectures. Additional results in Appendix B.2 are provided to confirm the previous insight. Surprisingly, Fig. 3(c) shows that fully-supervised models are better able to predict most of the regions than self-supervised models. Appendix B.1 provides the results of the previous study conducted on BOLD moments dataset across 46 regions, which confirms the consistency of our main findings across different datasets. Moreover, Appendix B.3 shows the model subfamilies while focusing only on the video understanding models and excluding the single image models, in addition to comparing convolutional vs. transformer based models that are trained with full supervision only.
@@
Figure 3: Real experiments showing regression scores as Pearson’s correlation coefficient of model families on brain fMRI data. Comparison between: (a) image vs. video understanding models, (b) convolutional vs. transformer-based models and (c) fully supervised vs. self-supervised models. Statistical significance is shown in the bottom as ‘ns’ not significant, ‘,,*,**,***∗ , ∗ ∗ , ∗ ∗ ∗’ significant with p-values $<0.05,0.01,0.001$, resp. It shows video understanding models outperform single image ones, fully supervised outperform self supervised ones and convolutional models surpass transformer based ones in early-mid regions.
Figure 4: Fine-grained analysis of the video understanding models across the nine regions of the human visual cortex showing the Pearson’s correlation coefficient as the regression scores. (a) Single stream vs. two stream SlowFast architectures. (b) OmniMAE pre-trained in a self-supervised manner vs. TimeSformer and OmniMAE fine-tuned with full supervision. All models are based on ViT-B and trained on SSV2. (c) Comparison between six video understanding models. Statistical significance is shown on top of bar pairs as ‘ns’ not significant, ‘,,*,**,***∗ , ∗ ∗ , ∗ ∗ ∗’ significant with p-values $<0.05,0.01,0.001$, resp.
@@ 4.4 Fine-grained analysis
In this section, we conduct a fine-grained analysis that goes beyond families of models. We start with studying two stream vs. single stream architectures across three video understanding datasets. Figure 4(a) shows that the two stream architectures have better ability to model the visual cortex than single stream ones in the low level regions and are either better or on-par in the high level regions. We then discuss the self-supervised learning results that showed worse regression scores in comparison to full supervision. Towards this end, we investigate OmniMAE variants (i.e., self supervised and finetuned) and TimeSformer. Figure 4(b) shows that fully supervised models give better scores than the self-supervised ones across the nine regions. We leave the reasons behind this result as an open question for future research on why fully supervised models tend to align better with the cortical responses than self supervised ones. Furthermore, we investigate which models are better at predicting the visual cortex responses. Figure 4(c) shows that both SlowFast, a two-stream architecture, and MViT, a multiscale vision transformer, are the best in modelling the visual cortex across the nine regions. SlowFast which is a convolutional approach is comparable to MViT with difference that is statistically not significant across all regions. Moreover, we see self-supervised models (i.e., stMAE and OmniMAE) lagging behind fully supervised ones. Additional analysis is provided in Appendix B.4 and B.5.
4.5 Inter-intra region connectivity
In this section, we show the results of our improved neural encoding mechanism that builds upon the best video understanding models (MViT-B and SlowFast) while incorporating intra- and inter-region connectivity. Figure 5(a) shows the statistically significant enhancements in prediction accuracy in both MViT-B and SlowFast models after incorporating voxel-connectivity to their predictions. These results show that integrating the intra- and inter-connectivity across the visual regions is an important component for a better neural encoding. As an ablation study, we examine the performance enhancement in MViT-B in the case of intra-connectivity or inter-connectivity separately. Figure 5(b) shows that the full-connectivity (i.e., combining both) is either superior or on-par with intra-connectivity or inter-connectivity standalone. To better understand the directional connectivity between the regions, we analyze the average learned weights of each region as contributors to the MViT-B accuracy enhancement of each target visual region as shown in Figure 5(c). It shows the following: i) the effect of one region on another is not symmetric but directional, ii) early-mid regions (V1, V2, V3, and V4) are the highest contributors to the accuracy enhancement of other early-mid regions, and the same for late-regions (LOC, EBA, FFA, STS, and PPA), iii) V4 is contributing to both early-mid and late regions, and iv) the contributions of late regions on early regions (V1, V2) are stronger than contribution of early regions on late regions which could be attributed to the top-down influence of feedback-pathways in the visual cortex Gilbert & Li (2013).
Finally, we perform an ablation of the improvement from intra- and inter-region connectivity prior when using an image understanding model vs. a video understanding one. Figure 6(a) illustrates the effect of inter-intra region connectivity priors on a single image understanding model (ResNet-50), where we show that it improves over the baseline with no connectivity across most of the regions with statistical significance. More importantly, we study the gain from the connectivity priors with that single image understanding model (ResNet-50) vs. a video understanding model that learns dynamics (MViT). Figure 6(b) shows that the gain from MViT is on average higher than the ResNet-50 across all the regions. When looking upon the statistical significance of these results, we notice almost half of the regions show a significant gain from encoding dynamics over the single image understanding baseline. Hence, we show in these results that the dynamics encoding in video understanding models reinforces the connectivity priors. We refer to additional ablations that compares our learned connectivity to a random or an identity one in Appendix B.6, again confirming on the benefit of our learned connectivity.
V1V2V3V4LOCEBAFFASTSPPA0.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.4Visual Cortex RegionsRegression ScoreIntra-connectivityInter-connectivityFull-connectivity
Figure 5: (a) Comparison of base model accuracies of MViT-B-16x4 and SlowFast and their accuracies after incorporating the intra-region and inter-region voxel connectivity showing the Pearson’s correlation coefficient as the regression scores. It shows the superiority of the connectivity based models. (b) Comparison of performance enhancement by incorporating the intra-region and inter-region voxel connectivity together or each of them separately showing the Pearson’s correlation coefficient as the regression scores. It confirms on the need for combining both intra- and inter-region connectivity. (c) Average weights per region contributing to the accuracy enhancement of each target visual region, showing the directional learned connectivity in our model.
Figure 6: (a) Comparison of base model accuracy of ResNet-50 and its accuracy after incorporating the inter-intra region voxel connectivity showing the Pearson’s correlation coefficient as the regression scores. (b) Comparison of improvement in the regression scores in ResNet-50 vs. MViT w/ connectivity priors. Difference in regression scores (i.e., Pearson correlation coefficient) is shown in a different scale after multiplying by 100 for visualization. It confirms that the dynamics based encoding improves the benefit from connectivity priors.
4.6 Summary
In summary: (i) Simulated results show that system identification between image and video understanding models is attainable to a certain level. (ii) We show that video understanding models are better in regressing the visual cortex responses than image ones. (iii) We show that convolutional models predict better the early-mid regions than transformer based ones. (iv) We show that MViT, with its multiscale processing, tends to perform similar to convolutional models in early-mid regions. (v) We show that two-stream models perform better than single stream. (vi) We show that models trained with full supervision surpass the ones with self supervision. (vii) Finally, we demonstrate a better neural encoding scheme that utilizes both encoded dynamics and inter-intra region connectivity. We refer to Appendix C and D for a discussion on the limitations and broader impact.
5 Conclusion
This paper has provided a large-scale study of video understanding models from a neuroscience perspective. We have shown the feasibility of system identification for single image vs. video understanding models. Moreover, we have shown that modelling dynamics should be considered when comparing biological neural system responses to deep networks and provided insights on different families of models. Finally, we showcased the interplay of dynamics modelling and inter-intra region connectivity in neural encoding.
6 Acknowledgements
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) and Digital Research Alliance of Canada (RAC) in conducting this research work.
Appendix A Candidate models details
In this section, we provide details on the layers extracted from each of our candidate models in Table 2. Note that for transformer-based architectures instead of using all layers to be selected we rather group layers into blocks of four layers for efficiency reasons and we noticed it gave better results than learning the regression with all input layers at once. Also note that OmniMAE finetuned is trained for the action recognition task on SSV2 dataset Goyal et al. (2017).
Table 2: Detailed description per architecture of the used layers.
Architecture # Layers Sampled Layers
CSN, R(2+1)D, 3DResNet 6 5 convolutional blocks and last fully connected layer
C2D, I3D 7 5 convolutional blocks, max pooling and last fully connected layer
SlowFast 12 5 convolutional blocks for each branch (slow & fast) and the last 2 layers combined
X3D 6 5 convolutional blocks and last fully connected layer
ViT, TimeSformer 4 Grouped 3 blocks, each block 4 layers and last fully connected layer
Dino - B 4 3 Grouped blocks (4 layers) and CLS token
ResNet 7 5 convolutional blocks, average pooling and last fully connected layer
MViT 5 Grouped 4 blocks (4 layers) and last fully connected layer
OmniMAE, MAE 3 Grouped 3 blocks (4 layers)
stMAE 6 Grouped 6 blocks (4 layers)
Appendix B Additional results
B.1 BOLD Moments Dataset (BMD) results
In this section, we evaluate on the newly released BOLD Moments Dataset (BMD) Lahner et al. (2024), with fMRI recording of ten subjects watching 1,000 training video stimuli. The dataset provides a preprocessed version (named “Version B”) with additional flexibility in the region of interest (ROIs) format. In “Version B” of BMD, 46 regions are defined including right/left hemispheres and ventral/dorsal streams. The defined ROIs include: early visual regions (V1v, V1d, V2v, V2d, V3v, V3d, hV4), a body-selective region (EBA), an object-selective region (LOC), face-selective regions (FFA, OFA, STS), scene-selective regions (PPA, RSC, TOS), additional ROIs (V3ab, IPS0, IPS1-2-3, 7AL, BA2, PFt, and PFop), and finally the motion selective MT ROI. All these ROIs were defined in both right and left hemispheres forming in total 46 ROIs. To further show the robustness of our results and insights, we evaluate all our models using this preprocessed version across the 46 ROIs using cross-validation on 4 folds as shown in Fig. 7. It confirms the consistency of the results previously shown in Fig. 3. Figure 7(a) confirms the statistically significant superiority of video understanding models compared to single-image models in the majority of cortical regions. Figure 7(b) confirms that convolutional-based models are better than transformer-based models in representing early-mid visual cortex regions. Finally, Fig. 7(c) confirms that fully-supervised models perform better than self-supervised ones in predicting activity of most of the visual cortex regions with statistical significance.
Figure 7: Real experiments showing regression scores as Pearson’s correlation coefficient of model families on brain fMRI data using the BMD dataset. (a) Comparison of image vs. video understanding models, (b) comparison of convolutional vs. transformer-based models and (c) comparison of fully supervised vs. self-supervised models. Statistical significance is shown in the bottom as ‘ns’ not significant, ‘,,*,**,***∗ , ∗ ∗ , ∗ ∗ ∗’ significant with p-values $<0.05,0.01,0.001$, resp.
B.2 Additional simulated experimental results
Although our focus in the system identification is on the ability to differentiate image vs. video understanding models, we provide additional results for other families of models. In Fig. 8, we show the simulated results comparing convolutional vs. transformer based models for four target models MViT-B 16x4, ViT-B, ResNet-50, and I3D ResNet-50, respectively. For the target model I3D, the figure clearly shows that convolutional models are better predictors than transformer-based ones with statistical significance across blocks except the last one. The target model ResNet50 shows that convolutional models are significantly better than transformer-based ones in the early blocks, but the differences are not significant in the late blocks. For ViT-B, the transformer-based models are significantly better than the convolutional models in all the blocks except for the first block which has non-significant difference between the two families. This result is consistent with the results presented in Han et al. (2023). The target model MViT-B shows a surprising result, that is confirming with previous findings in the real experiments as well as detailed in Section B.4, where convolutional models are better in regressing the multiscale variant than transformer-based ones with significant differences in the first two blocks. This might be related to how the multiscale ViT tends to act similar to the convolutional models when predicting early-mid regions of the visual cortex.
In the MViT target model results showed in Fig. 2(a) and Fig. 8(a), we did not include the MViT-32x3 variant for fair comparison between model families given that MViT-32x3 share the same architecture as MViT-16x4. However, for further investigation and understanding of MViT target model results, we re-assessed the results after adding MViT-32x3 as a source model. Figure 9 shows the results of MViT-16x4 target model after including MViT-32x3 source model. Figure 9(a) shows consistent results as shown in Figure 2(a) in which video understanding models are significantly better than image understanding ones. Figure 9(b) also shows consistent results to Figure 8(a) in which convolutional models are better than transformer-based models as predictors of MViT target model but with insignificant results, after adding MViT-32x3 source model to the transformers family, as anticipated.
Figure 8: Simulated experiments showing regression scores as Pearson’s correlation coefficient of (a-d) convolutional (blue) vs. transformer (red) model families on four target models MViT-B 16x4, ViT-B, ResNet-50, and I3D ResNet-50, respectively. Statistical significance is shown at the bottom as ‘ns’ not significant, ‘,,*,**,***∗ , ∗ ∗ , ∗ ∗ ∗’ significant with p-values $<0.05,0.01,0.001$, respectively.
Figure 9: Simulated experiments for MViT-B 16x4 target model after including MViT-B 32x3. Statistical significance is shown at the bottom as ‘ns’ not significant, ‘,,*,**,***∗ , ∗ ∗ , ∗ ∗ ∗’ significant with p-values $<0.05,0.01,0.001$, respectively.
B.3 Additional real experimental results
We add a study of the model subfamilies in terms of (a) convolutional vs. transformer based models and (b) fully vs. self supervised models, but focused on video understanding models only excluding models trained with single images. As shown in Fig. 10(a) it shows consistently that convolutional based models perform better in early layers as found earlier, where the first three regions show statistically significant results. Moreover, Fig. 10(b) shows models trained with self-supervision tend to be worse than fully supervised ones. However, note that these results are using a relatively small number of self-supervised learning video understanding models. Thus, we leave it for future work to expand on this further. Finally, we study the convolutional vs. transformer based models with focus on fully supervised ones excluding self supervision. Figure 10(c) shows the consistency of our results where the convolutional models in the early-mid regions surpass the transformer based ones.
Figure 10: Real experiments showing regression scores as Pearson’s correlation coefficient of model families on brain fMRI data with focus on video understanding models. (a) Comparison of convolutional vs. transformer-based amongst video understanding models. (b) Comparison of fully vs. self supervised amongst video understanding models. (c) Convolutional vs. transformer based fully-supervised models. Statistical significance is shown in the bottom as ‘ns’ not significant, ‘,,*,**,***∗ , ∗ ∗ , ∗ ∗ ∗’ significant with p-values $<0.05,0.01,0.001$, resp.